I have a word problem that I would like to check to see if i got it right. Please check my answer if you would.
Carlos can paint a house in 15 days. Carlos and Michael working together can do the same job in 10 days. How long would it take michael to do the job working alone? Answer to the nearest tenth of a day.
I got 10.0 days.
Thanks for checking.

Answer by Nate(3500)

Let's Check Your Answer:
Carlos' rate: 1/15
Both's rate: 1/10
Rate * Time = Work
t/10 - t/15 = 1
30t/10 - 30t/15 = 30
3t - 2t = 30
t = 30
Micheal can paint a house in 30 days.